OpenStudy (anonymous):

A train travels 20 feet in 1/5 second. At this same speed, how many feet will it travel in three seconds?

OpenStudy (anonymous):

if it covers 20 ft in 1/5 of a second then in 1 full second it will travel 20*5 = 100 ft so in three seconds it will travel = 100 * 3 = 300 ft

OpenStudy (anonymous):

Thank you so much!!!!! Awesome!!

OpenStudy (anonymous):

u r welcome ( :)>